OpenStudy (anonymous):

(2/3)^abslx+2l=(81/16)^-3

OpenStudy (anonymous):

\[(\frac{81}{16})^{-3}=(\frac{16}{81})^3=(\frac{2^4}{3^4})^3\] \[=(\frac{2}{3})^{12}\]

OpenStudy (anonymous):

this tells you \[|x+2|=12\] so \[x=2=12\] or \[x+2=-12\]